OpenStudy (anonymous):

The volume of a fudge tin meust be 120 cm^3. The tin is 7cm longer than it is wide and six times longer than it is tall. Find the dimenmsions of the tin

OpenStudy (anonymous):

l = w + 7 h = 6 * l = 6 * (w + 7) Area = w * l * h = 120

OpenStudy (calculusfunctions):

Alright, so here the shape of the tin would be a rectangular prism. Let w represent the width, (w + 7) represent the length, and then the height would be the length divided by 6 because the tin is six times longer than it is wide. Now the volume is 120 cm³, thus\[w(w +7)\left( \frac{ w +7 }{ 6 } \right)=120\]because the volume of a rectangular prism is the product of the area of the base and the height. Sorry it took so long but my server always gives me problems!!!

OpenStudy (calculusfunctions):

@kris10 do you understand the equation and do you know how to solve it now?

OpenStudy (anonymous):

so i'll just find the value of w and calculate each

OpenStudy (anonymous):

isn't it (6)(wt7)

OpenStudy (calculusfunctions):

yes! Why don't you tell me what the value of w is here so that I know you did it correctly.

OpenStudy (calculusfunctions):

No to (6)(wt7)!

OpenStudy (calculusfunctions):

Think about it! The question states that ...it is six times longer than it is tall. This implies that the length is greater than the height. However, if the height is 6(w + 7) then that makes the tin taller than the length. So 6(w + 7) for the height is wrong!!! For example, suppose that the the height is 10, the the length would be (6)(10) because it is six times longer than it is tall. Now if the length is 60, then to get the height, you would divide the length by 6. Do you understand @kris10 ?

OpenStudy (anonymous):

ok

OpenStudy (calculusfunctions):

OK so what is your answer? What is w?

OpenStudy (calculusfunctions):

What are the dimensions? Length, width, and height?

OpenStudy (anonymous):

A=(w^3+14w^2+49)/(6)

OpenStudy (calculusfunctions):

It's not A for area!!! You mean to say V for volume! Where are you getting A from?

OpenStudy (anonymous):

ok

OpenStudy (calculusfunctions):

Also your expanded form is incorrect because the last term in that trinomial should be 49w, NOT 49.

OpenStudy (calculusfunctions):

Now do you know how to solve the equation?

OpenStudy (anonymous):

yeah i saw the difference

OpenStudy (calculusfunctions):

OK so did you get the correct answer then?

OpenStudy (anonymous):

[6(120)]-49=w(w+7)^2

OpenStudy (calculusfunctions):

What?? HUH?? Please try again!

OpenStudy (anonymous):

i'm confused on how to get the value of w

OpenStudy (calculusfunctions):

\[w(w +7)\left( \frac{ w +7 }{ 6 } \right)=120\]Multiply both sides by 6 to obtain\[w(w +7)(w +7)=720\]Expand the left side and subtract 720 from both sides to obtain\[w ^{3}+14w ^{2}+49w -720=0\]Now apply the factor theorem to solve for w. @kris10 can you take it from here?

OpenStudy (calculusfunctions):

Try positive factors of 720 for w, to see which one makes the left side zero.

OpenStudy (calculusfunctions):

For example, the equation does not equal zero if w equals 1, 2, 3, or 4. Now you try 5 because it is the next factor of 720.

OpenStudy (anonymous):

i don't know i'm really getting headache with this

OpenStudy (calculusfunctions):

So we have w³ + 14w² + 49w - 720 = 0 If w = 5, then 5³ + 14(5)² + 49(5) - 720 = 125 + 350 + 245 - 720 = 0 ∴ the width is 5 cm, the length is 12 cm, and the height is 2 cm. @kris10 do you understand?

OpenStudy (anonymous):

yeah that's the answer but how do you derived that 5

OpenStudy (anonymous):

where did you get that 5

OpenStudy (calculusfunctions):

I already explained above that you have to try positive factors of 720.

OpenStudy (calculusfunctions):

Have you never learned the factor theorem? You would have had to in order to be assigned this question.

OpenStudy (anonymous):

step by step in getting the 5 as the value of w

OpenStudy (anonymous):

ok thanks a lot thank you for helping

OpenStudy (calculusfunctions):

So you get it then?

OpenStudy (anonymous):

can u help me in the next question

OpenStudy (calculusfunctions):

Sure, as long as it doesn't take as long as this one. Only because I have to go, otherwise I wouldn't mind that you take all the time it takes to understand.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The clothes adds a 75% mark up to the wholesale price of merchandise to determine the selling price. If a jacket that costs $100 wholesale is on sale for 20% off, what could a customer pay for it?

OpenStudy (calculusfunctions):

OK so first you tell me what you think we should do because even though I'm in a rush, as a teacher, I can't just keep giving you the answers. So go ahead and tell me what you think we should do.

OpenStudy (anonymous):

the price of the jacket? i think

OpenStudy (calculusfunctions):

OK look, just relax. I'll walk you through. First we're told that there is a markup of 75% to the wholesale price to determine the selling price, right? So first we setup an equation that represents the selling price. Reread the first sentence of the problem and literally translate the English sentence into a mathematical statement (equation). What do you gather the equation is?

OpenStudy (anonymous):

price=$100*(75%+20%)

OpenStudy (calculusfunctions):

No, forget about the 20% right now. Just concentrate on the first sentence. What is the equation for the selling price? You can't do the entire question in a single step.

OpenStudy (anonymous):

Selling price=wholeprice+75%

OpenStudy (calculusfunctions):

Close! It's Selling Price = wholesale + 75% of wholesale Let P represent the selling price and W represent the wholesale price, the the equation should be P = W + 0.75W OR P = 1.75W Understand?! Convert the percent to decimal.

OpenStudy (calculusfunctions):

Now utilize the second sentence of the problem. So what do you think we should do now?

OpenStudy (anonymous):

my answer is to to it heheh yes. . . i understand

OpenStudy (anonymous):

selling price($100x20%)

OpenStudy (calculusfunctions):

No, remember that P = 1.75W Now the second sentence of the problem vocalizes that the jacket costs $100 wholesale and is on sale for 20% off so then we'll let the D represent the dicounted price, which is the price that the customer will pay. First if the wholesale price is $100, then the regular price P is P = 1.75(100) or $175 Now if the discount applied is 20% to the regular price, then D = 175 - 20% of 175 OR D = 175 - 0.20(175) D = 175 - 35 D = 140 ∴ the customer pays $140 for the jacket after the discount. Understood?

OpenStudy (anonymous):

thanks a lot

OpenStudy (calculusfunctions):

So you're good with it then?

OpenStudy (anonymous):

a little bit thank you so much SIR